OpenStudy (anonymous):

solve for x: x^3-10x^2+24x

OpenStudy (anonymous):

u can divide for x

OpenStudy (anonymous):

it will be x^2-10x+24

OpenStudy (anonymous):

but u must writing \[x \neq 0\]

OpenStudy (anonymous):

it need to be an equation, to get the value for x

OpenStudy (lexi724):

messaging you

OpenStudy (anonymous):

D = b^2 - 4ac

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@Ziferot to apply for quadratic formula you need to make the expression =0

OpenStudy (lexi724):

no problem @Jake_Litt and What? @lall

OpenStudy (anonymous):

@lall ok x^2-10x+24=0

OpenStudy (anonymous):

thats batter??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it doesnt matter

OpenStudy (anonymous):

do you know\[ \frac{ d }{ dx }(x ^{2}-10x+24)=\pm \sqrt{10^{2}-4(1)24}\] @Ziferot

OpenStudy (anonymous):

???

OpenStudy (anonymous):

new for you? i was just sharing knowledge just solve it

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i have a very good knowledge in math and high math , but bad knowledge in English )))

OpenStudy (anonymous):

you get x=6,4

OpenStudy (anonymous):

solving it by quadratic formula, you get the same result

OpenStudy (anonymous):

I haven't resolve it ))) i'm on work)))

OpenStudy (anonymous):

but you don't need the quadratic formula

OpenStudy (anonymous):

yes this can be done by splitting method also

OpenStudy (anonymous):

+ yeah,his right )

OpenStudy (anonymous):

see this @Ziferot \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

\[2ax+b=\pm \sqrt{b ^{2}-2ac}\]

OpenStudy (anonymous):

\[\frac{ d }{ dx }(ax ^{2}+bx+c)=\sqrt{D}\]

OpenStudy (anonymous):

@Ziferot

OpenStudy (lexi724):

he already has the answer guys

OpenStudy (anonymous):

have you seen it before????

OpenStudy (anonymous):

its not correct

OpenStudy (anonymous):

how

OpenStudy (anonymous):

see it again

OpenStudy (lexi724):

in his inbox its not posted here